9x - 6y = -12
x + 2y= 0
how do you solve the system of equations by elimination

Answer by persian52(161)

9x-6y=-12_x+2y=0
=►Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 6.
9x-6y=-12_3*(x+2y=0)
=►Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 6.
9x-6y=-12_3*(x+2y)=3(0)
=►Multiply 3 by each term inside the parentheses.
_3*(x+2y)=0
=►Multiply 3 by each term inside the parentheses.
_(3x+6y)=0
=►Remove the parentheses around the expression 3x+6y.
_3x+6y=0
=►Add the two equations together to eliminate y from the system.
3x+6y=0_9x-6y=-12_12x =-12
=►Divide each term in the equation by 12.
x=-1
=►Substitute the value found for x into the original equation to solve for y.
9(-1)-6y=-12
=►Multiply 9 by each term inside the parentheses.
-9-6y=-12
=►Move all terms not containing y to the right-hand side of the equation.
-6y=-3
=►Divide each term in the equation by -6.
y=(1)/(2)
=►This is the final solution to the independent system of equations.
x=-1_y=(1)/(2)